A machinist wishes to insert a silver rod with a diameter of 4 mm into a hole with a diameter of 3.993 mm. By how much would the machinist have to lower the temperature (in °C) of the rod to make it fit the hole?


1
Expert's answer
2020-08-04T16:00:25-0400

The change in the linear dimension of the rod due to temperature can be estimated to be:


"\dfrac{\Delta d}{d} = \alpha\Delta T"

where "d = 4mm" is a a diameter of the rod, "\Delta d = 3.993 mm-4mm = -0.007mm" is the diameter change, "\alpha = 18\times 10^{-6} \degree C^{-1}" is a coefficient of linear thermal expansion and "\Delta T" is temperature change.

Expressing "\Delta T", obtain:

"\Delta T = \dfrac{1}{\alpha}\dfrac{\Delta d}{d} = \dfrac{-0.007}{18\times 10^{-6}\cdot4} \approx -97.2\degree C"

Answer. The machinist have to lower the temperature by 97.2°C.